### v4

# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manusript

Abstract
Despite recent advancements in deep learning, its application in real-world medical settings, such as phonocardiogram (PCG) classification, remains limited. A significant barrier is the lack of high-quality annotated datasets, which hampers the development of robust, generalizable models that can perform well on newly collected, out-of-distribution (OOD) data. Self-Supervised Learning (SSL), particularly contrastive learning, has shown promise in mitigating the issue of data scarcity by leveraging unlabeled data to enhance model robustness and effectiveness. Even though SSL methods have been proposed and researched in other domains, works focusing on the impact of data augmentations on model robustness for PCG classification is limited. In particular, while augmentations are a key component in SSL, selecting the most suitable transformations during the training process is highly challenging and time-consuming. Improper augmentations can lead to substantial performance degradation, even hindering the network’s ability to learn meaningful representations. Addressing this gap, our research aims to explore and evaluate a wide range of audio-based augmentations and uncover combinations that enhance SSL model performance in PCG classification. We conduct a comprehensive comparative analysis across multiple datasets and downstream tasks, assessing the impact of various augmentations on model performance and generalization. Our findings reveal that depending on the training distribution, augmentation choice significantly influences model robustness, with fully-supervised models experiencing up to a 32% drop in effectiveness when applied to unseen data, while SSL models demonstrate greater resilience, losing only 10% or even improving in some cases. This study also sheds light on the most promising and appropriate augmentations for robust PCG signal processing, by calculating their effect size on model training. These insights equip researchers and practitioners with valuable guidelines for building more robust, reliable models in PCG signal processing.
Index Terms:
Contrastive learning, deep learning, OOD representation learning, phonocardiogram classification, self-supervised learning=-21pt
I Introduction
Abnormal phonocardiogram (PCG) signals are often regarded as an important indicator of heart diseases or deficiencies linked to increased mortality [1]. Since the collection of PCG signals is non-invasive, the automatic, accurate detection of abnormalities in a sample has the potential to assist in early prevention or timely treatment of such diseases, especially when access to healthcare services for certain populations is limited [2].
The efficacy of automated methods for detecting abnormal audio samples, depends to a large extent on the quality and nature of the audio recordings themselves. Typically, signals are first manually analyzed and processed before being used to train classifiers that discriminate between normal recordings and samples that include indicators for different diseases [3]. However, these classifiers often fail to produce generalizable and robust representations of signals due to distribution shifts present in either the training data or on the data that the model is evaluated on after training [4]. For example, a robust classifier should be able to maintain its accuracy on signals that have been collected in different environments, from multiple devices, and with varying sampling rates. Specifically for the case of PCG recordings, additional factors such as microphone placement [5, 6], body size, and skin type [7, 8], among others, can significantly affect the signals. The above cases prove even more challenging when screening an unborn child/fetus, where PCG can only be captured through the mother’s abdomen [9]. In addition to recording conditions, the training of accurate ML models is dependent on annotated data, the collection of which proves highly time- and cost-consuming. The above is especially important in the biomedical domain, where available samples are scarce to begin with. As a result, classification models that have been trained on data originating from a single source or on low-quality signals, fail to generalize to previously unseen data distributions, which do not adhere to the i.i.d. assumption (i.e., are not independently and identically distributed), leading to degraded effectiveness [10, 4].
A possible solution to the above limitations, which has proven advantageous in various ML fields, is the utilization of unlabeled data [11, 12]. By training models on a large corpus of unlabeled data via self-supervised learning (SSL), the feature extractor is able to learn general representations and avoid overfitting on its training distributions. Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on either electrocardiography, electromyography or electroencephalography signals (see Section II-B). Therefore, in this work we turn our attention to the Phonocardiogram domain and propose training robust classifiers for PCG signal classification via contrastive self-supervised learning, leveraging both labeled and unlabeled datasets. Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13]. The design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework. To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations () and models () applied on PCG signals and report their effect on training and generalization. Specifically, we attempt to provide insight into the following research questions:
Does the application of contrastive SSL prove a valid training method for extracting robust and meaningful representations in the PCG classification domain?
If so, which augmentations or transformations lead to such representations and prove the most effective, and which actually inhibit training?
Is a classifier trained on the above representations able to maintain adequate effectiveness and demonstrate domain generalization capabilities when evaluated on OOD data, in contrast to its fully supervised counterpart?
To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets; (b) examine the effectiveness on a global “normal vs. abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable); (c) evaluate all trained classifiers on OOD data that are left out during training; and finally, (d) calculate the effect size of each augmentation. To the best of our knowledge, this is the first paper to research the effect of each applied augmentation on learning robust OOD PCG representations. To facilitate further research in the field and provide a comprehensive baseline, we have open-sourced our codebase111Code available at: https://github.com/aristotelisballas/listen2yourheart.
The rest of the paper is structured as follows. Initially, Section II presents previous work on PCG-based classification, SSL training, and model robustness. Subsequently, Section III presents the methodology, including training approach, network architecture, augmentations, and datasets. Finally, Section IV presents the evaluation framework and results, while Section V concludes the paper.
II Related Work
Accurate and robust PCG classification has major implications in the well-being of patients as it can often be an indicator of heart deficiencies. However, the limited availability of annotated and quality data has hindered the progress of research in the field, along with biosignal classification in general. In this section, we focus on the most important papers and previous works in DL for PCG classification and in SSL for 1D biosignal processing. We also refer to methods which have been specifically developed for robust 1D biosignal classification.
II-A Deep Learning for Phonocardiogram Classification
The release of publicly available PCG datasets under the scope of the 2016 [14] and 2022 [15] PhysioNet [16] challenges, has a sparked a line of research in deep learning for PCG classification. The majority of previous works have mainly proposed either 2D or 1D CNN architectures for tackling this particular problem. The main difference lies in whether the proposed models are trained directly on the one dimensional recorded auscultation waves or on spectral images extracted from the audio recordings. For example, multiple recent papers [17, 18, 19, 20] propose combining mel-spectrogram images extracted from PCG signals with encoded patient demographic data, in a dual-branch CNN model. The authors find that the combination of the above representations leads to improved accuracy in the heart murmur detection task. Similarly, in another work utilizing frequency domain features, the authors of [21] design a hierarchical deep CNN architecture that extracts and combines features from multiple-scales of PCG Mel-Spectrograms. A more recent work, [22], proposes training an ensemble of 15 ResNets [23] with channel-wise attention mechanisms on similar spectrograms, for grading murmur intensity in PCG recordings.
Although few, 1D models also prove promising in the heart-murmur detection domain. In [24] the authors employ a transformer-based neural network and train it on 1D wavelet power features extracted from raw PCG signals. On the other hand, [25] proposes using a U-Net to predict heart states from the raw PCG signal and combine the predictions with a ResNet model for the final classification. In a different approach, the authors of [26] suggest that the combination of a 1D CNN-Bi-LSTM model coupled with attention mechanisms, trained in conjunction with a feature-based classifier leads to robust abnormal PCG signal detection.
In this work, we build upon our preliminary research [27] for the 2022 PhysioNet Challenge. In particular, we demonstrate the value of adopting self-supervised contrastive learning for robust PCG classification and also report the top augmentations that should be implemented in such a framework. Through our research, we aim to extensively investigate the benefits of each augmentation, and their combinations, for contrastive SSL in 1D PCG signal processing. The aim of this study is to provide future researchers with best practices regarding PCG signal augmentation, while also proposing a structured protocol for the evaluation of developed methods.
II-B Self-Supervised Learning in 1D biosignal classification
Self-supervised representation learning (SSL) [28] was initially proposed for mitigating the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks [29] to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding [30], the goal of Self-Supervised Contrastive Learning [31, 32] is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions [33]. The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which capable of leveraging small amounts of data while yielding adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [34]. The authors of [35] propose augmenting multi-channel EEG signals by recombining several channels into additional ones which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying predefined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of [36] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in [37] where contrastive predictive coding is implemented on similar SSL-learned features to predict sleep stages and detect pathologies in multi-channel EEG signals. Finally, [38] proposed BENDR, which combines attention mechanisms [39] and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
Several SSL methods have been proposed for ECG signal analysis as well [40], most of which focus on either emotion recognition or pathology detection. For example, in [41] the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the-art results. In another paper studying emotion recognition from ECG signals, [42] introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44] or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, [46] propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models [47]. For example, [48] introduces NeuroPose, a system for 3D hand pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
II-C Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [50]. To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data-generating processes. Here, we reference prior works which fall under the above research fields.
Domain adaptation (DA) methods are the most popular choice when aiming to produce models with improved generalization capabilities in 1D biosignal classification. Similar to Transfer Learning (TL), DA algorithms leverage the representation learning capabilities of models pretrained on diverse source data distributions and fine-tune them on previously unseen but similar target distributions. The end-result is a model which remains robust under the distribution shift between source and target data. In healthcare, DA methods are proposed for a plethora of biosignal classification tasks. In [55], the authors propose pretraining vanilla CNN networks on a large corpus of raw ECG signals and demonstrate their effectiveness when fine-tuned on a smaller dataset for arrhythmia detection. In another more recent work, He et al. [56] propose tackling the distribution shift between different arrhythmia detection ECG datasets, by employing three distribution alignment mechanisms. To be more specific, they first extract spatio-temporal features from preprocessed signals, pass them through a graph convolutional neural network and finally align the resulting representations between source and target domains. Many studies have researched the advantages of DA and TL in EEG classification as well [57]. To name a few, the authors of [58, 59] and [60] propose aligning source and target distributions via adversarial learning for motor image EEG classification and epilepsy detection respectively, while [61] investigates several state-of-the-art DA algorithms for EEG-based emotion recognition.
Finally, Domain Generalization (DG) algorithms aim to tackle model generalization at its core, as the main goal is to provide models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of [62] propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, [63] and [64] propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of [4] introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from [65], for 12-lead ECG and 64-channel EEG classification.
III Materials and Methods
In this work, we introduce an evaluation framework for assessing the combination of various transformations, as well as the robustness of a network trained via contrastive SSL, against a fully supervised baseline model. This allows us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain, while also highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
For the self-supervised training of all models, we adapt the SimCLR [66] methodology to 1D PCG representation learning and evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
In the following sections, we describe the datasets and preprocessing steps followed for all signals included in the study, provide a detailed description of the protocol for evaluating the significance of each combination of augmentations, as well as the generalization ability of a model pretrained with self-supervised contrastive learning. Finally, we present the architecture and relevant implementation details of the 1D CNN network implemented for PCG classification.
III-A Datasets & Downstream tasks
The PCG signals used in this study originate from a total of 5 distinct public datasets and are available as “wav” files. These include the Ephnogram [67], FPCGDB [68], Pascal [69], PhysioNet2016 [14], and PhysioNet2022 [15] databases. Here, we describe each dataset separately and provide all necessary details regarding the preprocessing steps followed to homogenize all signals into a common format.
The Ephnogram [67] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals making it appropriate for developing noise-resilient methods. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled PCG signals for the SSL pretraining of our models. All recordings were sampled at 8 kHz.
The Fetal PCG Database (FPCGDB) [68], as its name suggests, is a dataset comprising 26 20-minute fetal phonocardiographic signals collected from healthy women during the final months of their pregnancies. The signals were sampled at 333 Hz. Similar to the Ephnogram dataset, these recordings were gathered for the development of fetal PCG signal simulation and fetal heart rate extraction algorithms, rather than for classification purposes, and thus do not include labeled samples. Additionally, many recordings are affected by noise, originating either from the pregnant body, such as muscular movements or placental blood turbulence, or from the surrounding environment.
The Pascal [69] Heart Sound Classification Challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole222Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
The remaining two datasets were also introduced as publicly available challenges for PCG classification. Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart sound recordings as a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
Finally, the PhysioNet2022 Challenge [15] was designed for the development of algorithms that can detect heart murmur in PCG signals. Building on the 2016 challenge, the 2022 edition contains a diverse dataset of heart sound recordings collected from multiple auscultation locations. The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals recorded from one of the four heart valves or another prominent auscultation location. Each recording is classified into one of three categories: present (heart murmur detected), absent (no heart murmur detected), or unknown (e.g., due to noise or corrupted audio), rendering this a multi-class classification task. All recordings were sampled at a rate of 4 kHz.
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process to [27] for converting all signals into a common format. As a first step and to avoid transient noise artifacts, we discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart sound recording333In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage we assign pseudolabels to each window, as described in Section III-B. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can also be evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section III-D.
III-B Self-Supervised Contrastive Learning for PCG Representation Learning
The main goal of this study is to first of all verify the use of SSL contrastive learning, but also evaluate the importance of augmentation selection for robust PCG representation learning. The proposed framework consists of two separate tasks; a Pretraining pretext task and a Downstream task, as depicted in Figure 1.
Assume a 1D PCG signal sampled from , where is the length of the signal. The aim of the contrastive SSL pretext task is to learn robust representations of from unlabeled PCG signals, by maximizing the agreement between two different augmented versions of itself (positive pair) and simultaneously minimizing the agreement between entirely separate augmented windows. Essentially, a model is trained to predict whether two augmented signals are the result of different transformations applied on the original sample. Formally, let be a transformation or augmentation function drawn from a family of functions . During the pretraining phase, an audio window is augmented to two different versions and , after the application of two transformation functions and respectively. Therefore, an initial batch of sample windows results into an augmented batch of signals , where and for . As mentioned, each augmented pair of signals resulting from the same original window is considered a positive pair, e.g , while pairs of different original windows comprise negative pairs, e.g . Following its augmentation, the signal batch is passed through a neural network encoder model (Figure 2), where each signal is encoded into a representation . Subsequently, the resulting representations are flattened and passed through a fully connected linear layer and projected into 128-dimensional vectors , as proposed in [70].
For the contrastive prediction task, we employ the NT-Xent loss function introduced in [30] and implemented in the SimCLR framework [70]. Specifically, given a set which includes an augmented positive pair and , a model trained via contrastive SSL aims to identify in for a given . The contrastive loss for a positive pair is formulated as follows:
where is the initial batch size of the randomly sampled signals, is an indicator function and a temperature parameter. The function calculates the cosine similarity between the encoded vectors and is defined as:
The final loss is computed across all positive pairs in a mini-batch.
After the pretraining step has been completed (i.e., step 2 in Figure 1), the weights of the backbone encoder network are frozen. At this point, the model is expected to produce a robust representation . While keeping the weights of the encoder frozen, a classification head , with randomly initialized weights, is attached to the pretrained network. The layers of the classification head are then trained on labeled audio windows in order to predict labels for each downstream task. The classification layers are trained via fully supervised learning, with the categorical Cross-Entropy (CE) loss. Depending on the labels of each dataset in the downstream task, the CE loss is reduced to its binary form.
III-C Augmentations for PCG Representation Learning
Arguably, the most integral part of the proposed framework is the family of transformation functions , which is applied on each audio window, as the type and intensity of each function can drastically affect the learning of the encoder network. In order to conduct a rigorous evaluation of all augmentations included in the proposed contrastive SSL framework and also report the functions which should most likely be applied for robust PCG representation learning, we select to implement functions that either simulate factors that can be found in real world scenarios. For example, the first thing that comes to mind is the volume of noise polluting recordings during PCG collection. Additionally, the difference in sensor conductance can lead to different scales of the recorded signals. To this end, it would make sense to at least either add artificial noise to randomly scale each audio window during the pretraining task. Following the above rationale, in our study we select to implement and evaluate the effectiveness of a total of 6 different augmentations, each of which is described below:
Noise: Zero-mean Gaussian Noise is added to the PCG signal. Specifically, given noise , where is the mean and the standard deviation, the resulting signal is .
Cut-off Filters: Here either a high-pass or low-pass filter is applied to the signal, at a 250, 500 or 750 Hz cutoff.
Scale: The signal is randomly scaled by a factor of to yield .
Reverse: The signal sequence is reversed.
Inversion: As its name suggests, this transformation inverts the original signal by multiplying it by -1, i.e., .
Random Flip: This function takes as input the signal and concurrently applies the ‘reverse’ and ‘inversion’ augmentations based on a given probability () between 0.3, 0.5 and 0.7.
In the proposed framework, we also evaluate the composition of augmentations by applying one transformation after the other. Specifically, we consider 4 different augmentation cases: 0vs1, 1vs1, 1vs2 and 2vs2. In each case, the signal is either not transformed, transformed once or augmented by the application of 2 concurrent functions.
III-D Evaluation Protocol
The present study aims at providing future researchers with a roadmap for applying SSL in Phonocardiogram processing problems. To this end, we design a rigorous and comprehensive protocol for evaluating the importance of each applied augmentation in robust PCG classification. Each different augmentation combination is evaluated by repeating the experiment illustrated in Figure 1, with the exact same data splits for fair comparison.
In the first step of the framework, we create a pool of 5 sec unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and Fpcgdb. After the preprocessing step has been completed and all signals are homogenized in the common format described in Section III-A, they are randomly shuffled and split into training and validation splits (80% and 20% respectively). In the second step, the split signals are cut into batches. Each sample in the batch is then duplicated and augmented by applying either one or a combination of the functions described in Section III-C. The augmented samples are then passed through a Neural Network encoder, which is trained via the contrastive pretraining task (Step 3). After the SSL backbone model has been trained, its weights are frozen.
In Step 4, a classification head with learnable weights is attached to the pretrained encoder which is then trained on data from one of the labeled datasets (Pascal, PhysioNet2016 or PhysioNet2022). Specifically, each of these datasets is split into training, validation and test splits with a standard 70%-20%-10% ratio and the trained model is evaluated on the test split only. Since each model is evaluated on test data originating from the same distribution as its training data, we call this the ”In-Distribution” evaluation. Inspired by recent work in biosignal domain generalization [4], we also design an evaluation protocol for assessing the robustness of a trained classifier on out-of-distribution data, which are drawn from entirely different datasets. Since the data has been completely left-out during downstream training, the above process is followed for evaluating the generalization ability of a model. Specifically, we now also freeze the weights of the attached classification layers and test the model on data drawn from the remaining 2 labeled datasets. For example, if the network is trained for the Pascal downstream task, it is evaluated on both in-distribution data (Pascal) and on out-of-distribution data (PhysioNet2016 and PhysioNet2022). We call this step the out-of-distribution or OOD evaluation step. However, in order to evaluate the same model on each separate dataset, all data must share common labels. To that end, we implement two different downstream task types: ‘All’ and ‘Binary’. In the ‘All’ downstream task, the original labels of each dataset are assigned to each audio window, while in the ‘Binary’ task a signal is labeled as either ‘normal’ or ‘abnormal’, as described in Section III-A. In the case of the PhysioNet2016 dataset, the two tasks match as it introduces a binary classification problem in the first place. For a fair comparison, all test data splits remain the same across each evaluation scheme. Finally, to assess the pretrained backbone encoder (and therefore the applied augmentations) on all available data, we repeat Steps 4 and 5 with the remaining datasets in a round-robin fashion (step 6 of the proposed framework). The significance of each augmentation and transformation is evaluated based on: (a) their occurence frequency in the top performing models in both In-Distribution and OOD evaluation splits, and (b) their effect on model training, as calculated by Cohen’s d measure [71]. Additional information and details regarding augmentation evaluation are provided in Section IV-A.
To thoroughly evaluate the proposed method we also repeat the entire experiment by leaving data out during Step 1 and pretraining the model with fewer data. This is also a means to concretely evaluate the proposed method on true OOD data, as they are completely hidden throughout the training process, either pretraining or downstream training. The best augmentation combinations are selected based on the effectiveness of the downstream models on the ‘In-Distribution’ test split. The same models are then evaluated on the hidden ‘OOD data’.
III-E Implementation Details
In this section, we describe all the implementation details regarding the execution of our experiments. First of all, for the backbone encoder model we implement the same model used in [27] and [72]. The encoder model is a 5 layer convolutional neural network, the architecture of which is depicted in Figure 2 (left). Given a batch size of the input of the encoder model is a tensor of dimension .
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the flattened output of the last layer, with an original dimension of into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq. 1). The dimension of the projection was set to follow the one in the original SimCLR paper [70]. For the cosine similarity (eq. 2) we use a temperature of 0.1, as in [27], while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer [73] with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.
During the downstream task, the dense projection layer of the pretrained encoder is discarded and all convolutional weights are frozen. Followingly, a 3-layered classification head is appended to the pretrained network, as shown in Figure 2 (right). The dimensions of each layer are set to reflect the model implemented in [27], which was shown to yield promising results. Dropout layers are included between each of the dense layers of the classification head in order to avoid overfitting. The application of dropout between multiple fully connected layers has shown improve the generalizability and robustness of classifiers [74, 72]. The output size of the last decision layer depends on the downstream task type and the label size in each dataset. For the ‘Binary’ type and the PhysioNet2016 dataset, the output size is set to 1. For the ‘All’ type, the output size is set to 5 or 3, for the Pascal and PhysioNet2022 datasets respectively. In all cases, the classification head is trained via the appropriate cross-entropy loss and the Adam optimizer. Additionally, the learning rate is set to and the batch size is set to 32. Once again, we train the model for 100 epochs at most, using early-stopping with 20 epochs patience. The same CNN model is used for the fully-supervised baseline. All experiments are implemented in Tensorflow 2.12 [75] and trained on a SLURM [76] cluster, containing 440GB NVIDIA A100 GPU cards, split into 8 20GB virtual MIG devices.
IV Results and Discussion
IV-A Augmentation and Transformation Evaluation
| Augmentation | Cohen’s () |
| Random Flip (0.5) | 1.8012 |
| Cut-off (500, 450) (LP) | 1.3295 |
| Cut-off (750, 700) (LP) | 1.2839 |
| Cut-off (250, 200) (LP) | 1.2639 |
| Random Flip (0.3) | 0.5784 |
| Random Flip (0.7) | 0.4923 |
| Scale (1.0, 1.5) | 0.2068 |
| Noise (-0.01, 0.01) | 0.1159 |
| Inversion | -0.0008 |
| Reverse | -0.0039 |
| Scale (1.5, 2.0) | -0.1169 |
| Cut-off (500, 550) (HP) | -0.1149 |
| Cut-off (250, 300) (HP) | -0.2079 |
| Noise (-0.001, 0.001) | -0.2398 |
| Noise (-0.1, 0.1) | -0.3257 |
| Scale (0.5, 2.0) | -0.4016 |
| Cut-off (750, 800) (HP) | -0.4743 |
To avoid the combinatorial explosion poised by the evaluation of all possible augmentation combinations, we designed a protocol to limit the experiments to a feasible amount. As mentioned in Section III-C, there are a total of 7 (including not transforming) possible functions that can be applied to the two copies of the audio sample, resulting in 4 possible augmentation cases; 0vs1, 1vs1, 1vs2 and 2vs2. As a first step, we examined all possible combinations in the 0vs1 and 1vs1 cases without applying the same transformation twice. After evaluating the above we selected a subset of augmentations which led to the top performing models in the “In-Distribution” setting and continued experimenting with 1vs2 and 2vs2 augmentation schemes. The models with the best results are presented in Table II and are discussed in the next section.
The main aim of this study is to provide an evaluation of the most appropriate augmentations for PCG signal analysis. Due to the fact that by only presenting the combination of the augmentations leading to the best classification results could be misleading (since only the specific combination could pertain to said results), we present a plot depicting the frequency of augmentation occurrence in the top 75 experiments across all 3 downstream tasks and for both In-Distribution and OOD data. Specifically, we select the top 25 results for each task and present the occurrences for all 75 total experiments in Fig. 3. The evaluation results are quite interesting as several conclusions can be drawn. First of all, the application of low-pass filters seem to be dominant in both In-Distribution and OOD top experiments. This is a clear indicator that such filters should be applied towards increasing the robustness of a classifier. What is more, the application of uniform noise, random flip and scaling on the signal, also boosts the effectiveness of classifiers in all cases. Furthermore, we found that the best performing models in each of the 3 downstream tasks, had adopted the following augmentations during their pretraining:
Pascal: low pass filter (pass band 250Hz, stop band 300 Hz) and random flip (p=0.7) vs inversion and uniform noise (mean 0, std 0.1),
PhysioNet2016: low pass filter (pass band 250Hz, stop band 300 Hz) vs reverse
PhysioNet2022: low pass filter (pass band 250Hz, stop band 300 Hz) vs random flip (p=0.7) and random scaling with between 1.0 and 1.5
While the above augmentations agree, for the most part, with the distributions and number of occurrences in Fig. 3, they also support the findings in several previous contrastive SSL works [66, 77], in which stronger augmentations lead to improved results.
To further validate our results and to draw concrete conclusions regarding the importance of each applied transformation, we calculate Cohen’s [71] for each implemented function. Cohen’s provides a standardized measure of the effect size of each augmentation and allows the comparison of their magnitude. Given two groups and , Cohen’s is the difference in their means divided by the pooled standard deviation (). Specifically, Cohen’s is defined as:
where the pooled standard deviation is defined as:
and , are the variances for each group respectively. The consensus on the interpretation of Cohen’s d is generally that its sign means to either a negative or positive effect, while a magnitude of up to corresponds to a very small or very large effect, respectively. To calculate Cohen’s , we create two subgroups for each augmentation. Specifically, let be an augmentation under consideration. The first group () contains all experiments where was implemented, while the second group () consists of all experiments with the same set of augmentations but without . For example, say that the augmentation under consideration is Noise. One experiment to include in the first group would be the Noise vs Inversion experiment. Therefore the corresponding experiment which should be added to the second group would be None vs Inversion. The same process can be repeated for all augmentations and for all runs, including experiments, as specified in Sec. III-C. Accordingly, we calculate the effect of each augmentation based on the Out-of-Distribution Micro F1 scores for the PhysioNet2016 dataset, when trained on signals from PhysioNet2022. The magnitude of effect regarding each augmentation is presented in Table I.
Interestingly enough, the statistics presented in Fig. 3, seem to agree with the effect sizes of each augmentation (Table I). Based on Cohen’s measure, the low-pass filters have the most important positive effect on robust SSL PCG learning. In simple terms, this means that the addition or application of low-pass cut-off filters during the pretraining phase of the encoder plays a significant role () in learning robust representations. The above result agrees with the general consensus that important biosignal artifacts are present in low frequencies, whearas unrelated noise is attributed to higher frequencies [78, 79, 80]. Additionally, the application of Reverse and Inversion (i.e Random Flip) has an overall positive effect on training, as shown by the positive measure. What’s more, the balanced reversion of inversion of the signals (i.e Random Flip with a .5% chance), has the largest positive effect (), whereas the application of only one of the above flip augmentations seems to have approximately no effect on training. Similarly, but to a less degree, scaling the signals by a small factor, assists the model in extracting meaningful representations. On the other hand, the addition of noise, the application of high-pass filters and large scaling, generally appear to have a negative effect. The intuition behind the negative effect of noise, is that since the selected signals derive from an already “noisy” dataset, the additive noise renders them meaningless, while the application of high-pass filters removes important signal artifacts present in lower frequencies.
Even though low-pass filters seem to be have the largest postive effect in training, they have to be combined with further transformations for a SSL model to reach optimal performance, as already mentioned. In the next section we present numerical results for all downstream tasks and experiments.
IV-B Results
| In-Distribution | Out-of-Distribution | Out-of-Distribution | |||||||
| All | Binary | Binary | Binary | ||||||
| SSL Datasets | Model | Acc | F1 | Acc | F1 | Acc | F1 | Acc | F1 |
| Pascal | PhysioNet2016 | PhysioNet2022 | |||||||
| — | Baseline | 79.37 | 48.37 | 67.77 | 40.00 | 51.95 | 62.74 | 51.33 | 59.41 |
| E, F, Pa, P16 | SSL | 76.87 | 42.18 | 31.25 | 47.61 | 63.55 | 74.88 | 53.24 | 64.44 |
| E, F, Pa, P22 | SSL | 76.87 | 42.18 | 32.10 | 42.25 | 61.56 | 72.69 | 61.11 | 73.64 |
| E, F, P16, P22 | SSL-NoDs | 75.62 | 39.06 | 31.25 | 40.00 | 59.88 | 74.97 | 58.33 | 72.06 |
| all | SSL | 81.25 | 53.12 | 35.93 | 47.61 | 67.13 | 79.09 | 62.11 | 74.72 |
| PhysioNet2016 | Pascal | PhysioNet2022 | |||||||
| — | Baseline | — | — | 88.64 | 93.09 | 55.00 | 40.00 | 49.85 | 60.59 |
| E, F, Pa, P16 | SSL | — | — | 80.07 | 87.38 | 48.58 | 42.90 | 57.74 | 70.97 |
| E, F, Pa, P22 | SSL-NoDs | — | — | 74.97 | 85.69 | 52.55 | 42.61 | 58.99 | 71.59 |
| E, F, P16, P22 | SSL | — | — | 84.19 | 89.65 | 31.25 | 47.62 | 62.07 | 78.07 |
| all | SSL | — | — | 85.02 | 90.31 | 42.19 | 50.66 | 70.45 | 81.63 |
| PhysioNet2022 | Pascal | PhysioNet2016 | |||||||
| — | Baseline | 90.80 | 86.21 | 86.90 | 91.90 | 53.12 | 53.12 | 67.91 | 80.64 |
| E, F, Pa, P16 | SSL-NoDs | 87.94 | 81.91 | 82.49 | 89.71 | 48.29 | 47.09 | 70.45 | 81.63 |
| E, F, Pa, P22 | SSL | 88.55 | 82.83 | 82.83 | 89.96 | 51.42 | 48.64 | 71.07 | 82.94 |
| E, F, P16, P22 | SSL | 85.37 | 78.06 | 78.06 | 87.67 | 31.25 | 47.61 | 74.97 | 85.70 |
| all | SSL | 89.31 | 83.96 | 83.67 | 90.00 | 48.44 | 54.79 | 74.97 | 85.70 |
Table II presents the Accuracy and F1-Scores of the best performing models trained via contrastive SSL for all downstream tasks, including evaluation on OOD distribution data, as presented in Section III-D. We perform a total of 4 cycles of SSL experiments, evaluating all augmentation combinations in each cycle. In each experiment cycle, we either use all available datasets or leave one labeled dataset out at a time. The first column of Table II denotes the datasets used during SSL training. In the same table, the ‘Baseline’ model refers to the fully-supervised model trained solely on data from the downstream ‘In-Distribution’ dataset. Due to the variance in the OOD results for the baseline model, we present the average of 5 total runs. Finally, the ‘SSL-NoDs’ model refers to the model which has not been pretrained with data originating from the dataset of the downstream task. All results are rated based on the F1-Scores on the ‘In-Distribution’ test splits and the adopted augmentations for each case are the ones mentioned in Section IV-A. The same models are then evaluated on the test splits of the OOD datasets.
IV-C Discussion
At a first glance, the benefits of SSL pretraining admittedly do not seem very compelling. When evaluating the model on the test data of the In-Distribution data, the baseline fully-supervised model outperforms the SSL models in 2 out of 3 cases. When trained on the PhysioNet2016 and PhysioNet2022 datasets the baseline surpasses the best SSL model by 2.78% and 1.9% respectively. However, the same does not hold in the case of the Pascal dataset, where the SSL model is able to outperform the baseline by 4.75% when evaluated on ‘All’ labels and by 7.61% in the ‘Binary’ task. Since the Pascal dataset is considerably smaller than the other 2, and arguably the hardest among the datasets, the above results validate the effectiveness of pretraining a model on a large corpus of data, when available labeled data is not sufficient.
By diving further into the results, the true benefits of the proposed method begin to show. First of all, the drop in effectiveness between in-distribution and OOD data is apparent. Additionally, in all cases the baseline models appear to have overfit their training data distribution and fail to maintain their classification ability across distinct datasets. On the other hand, the models which have been pretrained via SSL yield far superior results. In all cases, the models trained with the proposed framework surpass the baseline by 11.68% on average and even outperform it by 21.04% in the case of the PhysioNet22 OOD task. Even more so, the SSL models were surprisingly able to surpass both the effectiveness of the baseline and their ‘In-Distribution’ counterpart, in the case of the Pascal OOD evaluation. Specifically, the downstream models trained on data from PhysioNet2016 and PhysioNet2022 were able to yield better results than the fully-supervised model trained solely on Pascal data. Furthermore, even though there is a decrease in effectiveness when the OOD datasets are left out during model pretraining, the SSL models continue to outshine the baseline across the board, with the exception of the case where the downstream task is the PhysioNet2022 data and models are evaluated on OOD data from Pascal. Finally, we would like to note that the above findings are aligned with the results of our preliminary SSL method [27] submitted in the PhysioNet2022 challenge, where the proposed model was able to avoid overfitting on the available training distribution and demonstrated adequate generalization effectiveness on completely hidden data.
To further demonstrate the generalization capabilities of the SSL models, we provide t-SNE visualizations and illustrate the feature distribution of representations extracted from the penultimate layer of the best-performing model, in addition to representations extracted from a fully-supervised baseline model. Specifically, we visualize the features from signals drawn from distributions the models have been trained on (In-Distribution) and unseen distributions (OOD) in Fig. 4. As depicted, the feature distribution of the baseline models are disordered, without a clear distinction between the abnormal (blue) and normal (signals). On the other hand, the features of the SSL model portray a somewhat clearer structure with the features of each signal type grouping into clusters towards the upper-right and bottom-left corners. The compelling aspect of the visualization results, lies in the fact that the general feature distributions of the SSL models align on both in-distribution and completely unseen during training OOD samples.
V Conclusions & Future Work
This work provides a thorough evaluation of augmentation techniques and transformations for pretraining neural networks to achieve robust phonocardiogram (PCG) classification. Our method focuses on identifying the most effective augmentations for improving the performance of self-supervised learning (SSL) models. We present both the top-performing combinations of augmentations and a summary of the most frequently used transformation functions in the top 75 experimental results. Our findings highlight that the use of low-pass cutoff filters, consistently leads to the best-performing models across various datasets, making them essential for this task. Additionally, adding small random noise and applying signal inversion or reversal further enhance model performance.
Our results validate that contrastive SSL significantly improves the robustness and generalizability of PCG classifiers. Compared to fully-supervised models, SSL models experience a much smaller drop in effectiveness when evaluated on unseen data. However, the proposed method has limitations: SSL pretraining requires access to a large data corpus to maximize the model’s representation learning capabilities, and the contrastive loss function relies on large batch sizes, imposing a computational constraint.
In future work, we plan to refine the augmentation techniques by introducing new functions and exploring additional parameters, extend our approach to transfer learning tasks, and investigate the most effective augmentation policies for other 1D biosignals such as ECG, EMG, and EEG classification.

### v5

@@
-Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manusript
+Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manuscript
 Abstract
@@
 By training models on a large corpus of unlabeled data via self-supervised learning (SSL), the feature extractor is able to learn general representations and avoid overfitting on its training distributions.
-Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on either electrocardiography, electromyography or electroencephalography signals (see Section II-B).
+Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on electrocardiography, electromyography or electroencephalography signals (see Section II-B).
 Therefore, in this work we turn our attention to the Phonocardiogram domain and propose training robust classifiers for PCG signal classification via contrastive self-supervised learning, leveraging both labeled and unlabeled datasets.
-Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13].
-The design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework.
-To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations () and models () applied on PCG signals and report their effect on training and generalization.
+Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13], the design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework.
+To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations (~) and models (~) applied on PCG signals and report their effect on training and generalization.
 Specifically, we attempt to provide insight into the following research questions:
-Does the application of contrastive SSL prove a valid training method for extracting robust and meaningful representations in the PCG classification domain?
+Does the application of contrastive SSL prove to be a valid training method for extracting robust and meaningful representations in the PCG classification domain?
 If so, which augmentations or transformations lead to such representations and prove the most effective, and which actually inhibit training?
 Is a classifier trained on the above representations able to maintain adequate effectiveness and demonstrate domain generalization capabilities when evaluated on OOD data, in contrast to its fully supervised counterpart?
-To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets; (b) examine the effectiveness on a global “normal vs.
-abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable); (c) evaluate all trained classifiers on OOD data that are left out during training; and finally, (d) calculate the effect size of each augmentation.
+To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets, (b) examine the effectiveness on a global “normal vs.
+abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable), (c) evaluate all trained classifiers on OOD data that are left out during training, and finally, (d) calculate the effect size of each augmentation.
 To the best of our knowledge, this is the first paper to research the effect of each applied augmentation on learning robust OOD PCG representations.
@@
 The SSL paradigm seems to fit quite well in the medical and healthcare domain.
-In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which capable of leveraging small amounts of data while yielding adequate and trustworthy inference results.
-To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
+In a field where sufficient annotated datasets are scarcely available, there is a need to provide systems which are capable of leveraging small amounts of data while yielding adequate and trustworthy inference results.
+To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG), or electromyography (EMG) signals.
 Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [34].
 The authors of [35] propose augmenting multi-channel EEG signals by recombining several channels into additional ones which represent different aspects of the EEG recording.
-On top of the combination, they augment each channel by applying predefined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification.
+On top of the combination, they augment each channel by applying predefined transformations and employ contrastive loss for improved emotion recognition and sleep-stage classification.
 In another work, the authors of [36] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task.
@@
 They evaluate their method on four emotion recognition datasets and report state-of-the-art results.
-In another paper studying emotion recognition from ECG signals, [42] introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals.
+Another paper [42] studying emotion recognition from ECG signals introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals.
 Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia.
 Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice.
-Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44] or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection.
+Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44], or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection.
 In a different approach, [46] propose learning good representations by contrasting signals both at the inter- and intra-subject levels.
 Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
-Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models [47].
-For example, [48] introduces NeuroPose, a system for 3D hand pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user.
-Similarly, the authors [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
+Finally, EMG signal processing is another interesting field that can greatly benefit from the application of robust prediction models [47].
+For example, [48] introduces NeuroPose, a system for 3D hand-pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a another user.
+Similarly, the authors of [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
 II-C Model Robustness in 1D biosignal classification
 Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [50].
-To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-i.i.d.
-distributions or out-of-distribution (OOD) data for model evaluation.
+To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-IID distributions or out-of-distribution (OOD) data for model evaluation.
 These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data-generating processes.
@@
 Although limited, there has been a recent interest in biosignal DG research.
-In the field of PCG classification, the authors of [62] propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification.
+In the field of PCG classification, the authors of [62] propose an ensemble classifier fusion method that yields improved results in a DG setup for abnormal vs normal signal classification.
 To tackle the domain shift and variability present in ECG signals collected from different patients, [63] and [64] propose domain-adversarial model training for ECG and EEG classification, respectively.
@@
 In this work, we introduce an evaluation framework for assessing the combination of various transformations, as well as the robustness of a network trained via contrastive SSL, against a fully supervised baseline model.
-This allows us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain, while also highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
+This enables us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain as well as highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
 For the self-supervised training of all models, we adapt the SimCLR [66] methodology to 1D PCG representation learning and evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
@@
 The Ephnogram [67] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions.
-Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds.
-As specified by the researchers, the dataset includes poor sample signals making it appropriate for developing noise-resilient methods.
+Most recordings (62 during exercise) are 30 minutes long while the duration of the remaining 7 (rest) is about 30 seconds.
+As specified by the researchers, the dataset includes poor-quality sample signals making it appropriate for developing noise-resilient methods.
 As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples.
@@
 The Pascal [69] Heart Sound Classification Challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies.
-The challenges contains two tasks; one for heart sound segmentation and one for classification.
+The challenge contains two tasks; one for heart sound segmentation and one for classification.
 In our work, we use the publicly available labeled data for model training and evaluation which was gathered in clinical and in-the-wild settings.
@@
 The remaining two datasets were also introduced as publicly available challenges for PCG classification.
-Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart sound recordings as a binary classification problem.
+Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart-sound recordings as a binary classification problem.
 The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds.
@@
 Building on the 2016 challenge, the 2022 edition contains a diverse dataset of heart sound recordings collected from multiple auscultation locations.
-The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals recorded from one of the four heart valves or another prominent auscultation location.
+The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals, recorded from one of the four heart valves or another prominent auscultation location.
 Each recording is classified into one of three categories: present (heart murmur detected), absent (no heart murmur detected), or unknown (e.g., due to noise or corrupted audio), rendering this a multi-class classification task.
@@
 As a first step and to avoid transient noise artifacts, we discard the first and final 2 seconds of each PCG recording.
-The resulting signal is split into 5 second overlapping windows with a 50% overlap.
-After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart sound recording333In this study, all recordings initially labeled as unknown are considered abnormal..
+The resulting signal is split into 5-second, overlapping windows with 50% overlap.
+After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart-sound recording333In this study, all recordings initially labeled as unknown are considered abnormal..
 For the SSL pretraining stage we assign pseudolabels to each window, as described in Section III-B.
@@
 For the contrastive prediction task, we employ the NT-Xent loss function introduced in [30] and implemented in the SimCLR framework [70].
-Specifically, given a set which includes an augmented positive pair and , a model trained via contrastive SSL aims to identify in for a given .
-The contrastive loss for a positive pair is formulated as follows:
-where is the initial batch size of the randomly sampled signals, is an indicator function and a temperature parameter.
+Specifically, given a set which includes an augmented sample pair and , a model trained via contrastive SSL aims to identify in for a given .
+The contrastive loss applied on the representations of any pair is formulated as follows:
+where and are the embedded representations of and respectively, is the number of the initially randomly sampled signals (before the augmentations, i.e., ), is an indicator function and a temperature parameter.
 The function calculates the cosine similarity between the encoded vectors and is defined as:
-The final loss is computed across all positive pairs in a mini-batch.
+The final loss is computed across all pairs in a mini-batch.
 After the pretraining step has been completed (i.e., step 2 in Figure 1), the weights of the backbone encoder network are frozen.
@@
 Arguably, the most integral part of the proposed framework is the family of transformation functions , which is applied on each audio window, as the type and intensity of each function can drastically affect the learning of the encoder network.
-In order to conduct a rigorous evaluation of all augmentations included in the proposed contrastive SSL framework and also report the functions which should most likely be applied for robust PCG representation learning, we select to implement functions that either simulate factors that can be found in real world scenarios.
-For example, the first thing that comes to mind is the volume of noise polluting recordings during PCG collection.
+In order to conduct a rigorous evaluation of all augmentations included in the proposed contrastive SSL framework and also report the functions which should most likely be applied for robust PCG representation learning, we select to implement functions that either simulate factors that can be found in real world scenarios, such as the level (volume) of noise.
 Additionally, the difference in sensor conductance can lead to different scales of the recorded signals.
@@
 Scale: The signal is randomly scaled by a factor of to yield .
-Reverse: The signal sequence is reversed.
+Reverse: The signal sequence is reversed, i.e., where is the length of signal .
 Inversion: As its name suggests, this transformation inverts the original signal by multiplying it by -1, i.e., .
@@
 Each different augmentation combination is evaluated by repeating the experiment illustrated in Figure 1, with the exact same data splits for fair comparison.
-In the first step of the framework, we create a pool of 5 sec unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and Fpcgdb.
+In the first step of the framework, we create a pool of 5-second, unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and FPCGDB.
 After the preprocessing step has been completed and all signals are homogenized in the common format described in Section III-A, they are randomly shuffled and split into training and validation splits (80% and 20% respectively).
@@
 Given a batch size of the input of the encoder model is a tensor of dimension .
-For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the flattened output of the last layer, with an original dimension of into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq.
+For the contrastive SSL pretraining phase, we train the convolutional layers of the network by projecting the flattened output of the last layer, with an original dimension of into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq.
 1).